OpenStudy (sh3lsh):

Determine the slope of y=2x^2-4x+2.

OpenStudy (inkyvoyd):

Slope>?

OpenStudy (inkyvoyd):

Do you mean the derivative?

OpenStudy (anonymous):

In this case, slope is gradient. this means you need to differentiate to make the quadratic equation into linear equation. \[y=2x ^{2}-4x+2\] \[dy/dy =4x-4\] the slope is 4. Hope this helps. Correct me if i'm wrong.

OpenStudy (sh3lsh):

You're great yukitou. We haven't even remotely gone over derivatives. Could you tell me how you made the 2x^2 go away?

OpenStudy (inkyvoyd):

@sh3lsh , are you looking to find the solution using the difference quotient, or have you already learned the rule \(\large \frac{d}{dx} ax^n=nax^{n-1}\)

OpenStudy (sh3lsh):

difference quotient.

OpenStudy (inkyvoyd):

Alright. Just for the record, the slope is not 4. But, I'll explain that later.

OpenStudy (inkyvoyd):

We shall take the difference quotient. Give me a moment to type it for this quadratic function.

OpenStudy (inkyvoyd):

\(\Huge \frac{f(x+\Delta x)-f(x)}{\Delta x}\) First, I must make sure you know what this difference quotient even represents. Do you know?

OpenStudy (sh3lsh):

Yes I do.

OpenStudy (inkyvoyd):

\(\huge \lim_{\Delta x->0}\frac{(2(x+\Delta x)^2-4(x+\Delta x)+2)-(2(x)^2-4(x)+2)}{\Delta x}.\)

OpenStudy (inkyvoyd):

Oh great. let me retype that -.-

OpenStudy (sh3lsh):

hahaha! No no! I understand the point!

OpenStudy (inkyvoyd):

Anyways, I'm going to multiply this out lik a fool, because I completely forget how I used to evaluate the difference quotient with limits.

OpenStudy (sh3lsh):

Haha, we did this in trig a bunch. I didn't know that derivative were tied in. You do NOT have to multiply it out!

OpenStudy (inkyvoyd):

difference quotient of 2x^2-4x+2 \(\huge \frac{2(x^2+2\Delta x+\Delta x^2)-4x-4\Delta x-2x^2+4x-2}{\Delta x}\)

OpenStudy (inkyvoyd):

difference quotient of 2x^2-4x+2 \(\huge \frac{2x^2+4\Delta x+2\Delta x^2-4x-4\Delta x-2x^2+4x-2}{\Delta x}\)

OpenStudy (sh3lsh):

and the any term with h at the end is just thrown away right? I got 4x-4 :)

OpenStudy (inkyvoyd):

Gah, it's 4x-4

OpenStudy (inkyvoyd):

this is why we have the d/dx ax^n=ax^(n-1)

OpenStudy (inkyvoyd):

and d/dx (u+v)=d/dx u+d/dx v

OpenStudy (inkyvoyd):

I hate recreating the wheel lol

OpenStudy (sh3lsh):

Haha, you helped a lot. Thanks again. So much.

OpenStudy (inkyvoyd):

Trust me, it's easier to just take the derivative in the future.

OpenStudy (anonymous):

Sorry for the late reply. How I made the 2x^2 go away? Actually this is pretty straight forward after you do a lot of differentiation question. \[y=2x ^{2}-4x+2\] From this equation, you know you want to differentiate the equation. So you put the power 2 in front of the 2x.  Do you now why "2" of this y=2x^2 -4x=2 go?